# Break the Checkbox: Challenging Closed-Style Evaluations of Cultural Alignment in LLMs

### v2

@@ 1 Introduction
To fully unlock the potential of artificial intelligence and achieve the vision of “AI for All”, it is essential to design and develop the Large Language Models (LLMs) with a focus on inclusivity and relevance across diverse societies and cultures (Adilazuarda et al., 2024). However, the recent advancements in language models have faced significant criticism for predominantly reflecting Western perspectives (Durmus et al., 2023) and displaying pronounced biases toward Anglocentric or American cultural norms (Johnson et al., 2022; Dwivedi et al., 2023). Such biases pose serious risks, including the stereotyping and alienation of users from underrepresented communities due to a lack of cultural sensitivity and understanding (Cao et al., 2022). Addressing this issue begins with identifying and understanding the problem. This raises a critical question: to what extent do current LLMs align with the diverse cultural contexts worldwide, and how can this alignment be systematically evaluated?
@@
  • We demonstrate that closed-style questions alone are insufficient for comprehensively evaluating LLMs for cultural alignment, as unconstrained prompts often yield richer and more insightful responses and exhibit stronger alignment across cultures.
@@ 2.1 World Values Survey
The World Values Survey (WVS) (Haerpfer et al., 2024) is a global research initiative consisting of approximately 250 questions, organized into 14 thematic sections covering topics such as social values, stereotypes, societal well-being, and economic values. Political scientists Ronald Inglehart and Christian Welzel analyzed WVS data and identified two key dimensions of cross-cultural variation: Traditional versus Secular-rational values and Survival versus Self-expression values. These dimensions are derived through a factor analysis of $10$ indicators presented in Table 5. Though these indicators represent only a subset of beliefs and values, they effectively capture critical aspects of cultural variation. We adopt these $10$ closed-style indicators/questions as our first probing corpus.
@@ 2.3 Probing Methods
We concatenate Anthropological Prompting (AlKhamissi et al., 2024) before each probing, which grounds questions in anthropological contexts by guiding the model to think as if actively participating in the method.The authors identified six demographic dimensions and empirically found that optimal cultural alignment occurs with the following specific attributes: Region: Country-Specific, Sex: Male, Age: $<50$, Social Class: Upper/Lower Middle Class, Education Level: Higher, and Marital Status: Married.We adopt this configuration to allow our models the best opportunity to align with the cultural contexts under study. Our complete probing methodology is illustrated in Figure 2, using one example question from both the WVS and Hofstede surveys.
@@
2.4 Country & Langugage Selection
To ensure a robust analysis of cross-cultural differences, we select countries based on their positions on the Inglehart-Welzel World Cultural Map 222www.worldvaluessurvey.org/WVSNewsShow.jsp?ID=467, which is structured around two key dimensions: Traditional versus Secular-rational values and Survival versus Self-expression values. This method allows us to capture contrasting cultural orientations while ensuring the availability of ground truth data for validation. Additionally, we prompt language models in both English and the native languages of the selected countries to enhance cultural alignment and correlate results with survey findings (Cao et al., 2023a). Based on these criteria, we select Bangladesh, Germany, and the USA. Bangladesh represents a society deeply rooted in Traditional and Survival values, while Germany aligns strongly with Secular-rational and Self-expression values. To provide a reference point, we include the USA, as it is frequently used in cross-cultural studies and because large language models (LLMs) often exhibit bias toward American culture (Johnson et al., 2022). The USA also occupies the Secular-rational and Self-expression dimension, similar to Germany on the Inglehart-Welzel map. For Bangladesh and Germany, we source survey questions in Bengali and German from the official survey websites, which provide translations in numerous languages. We assume English to be the native language of the USA in our study.
@@ 3.1 Models
Table 1: Cultural alignment comparison using World Values Survey (WVS) when prompted in English and Native languages for Germany, Bangladesh, and the USA. Scores are in Hard / Soft alignment metrics for each model using four probing method: FC (Forced Closed-Style), FR (Forced Reverse Order), FO (Forced Open-ended), FU (Fully Unconstrained). Boldfaced are the highest alignment scores per country across the four probing methods for each model. The columns for the USA remain the same across languages assuming English to be the native language.
@@ 3.2 Evaluation
For the first two prompting methods, forced closed-style and forced reverse order, model responses are obtained in numerical form, consistent with traditional survey options. We use a straightforward Python script to map the responses from the reverse order questions back to the original survey option for evaluation. For the unconstrained settings (forced open-ended and fully unconstrained), we follow the approach of Bravansky et al. (2025) to map free-form responses to the most appropriate survey options. Specifically, we use GPT-4o to analyze these responses and determine the model’s stance using the prompt shown in Figure 6. In many instances, LLMs respond to certain questions with statements like, “As an AI, I do not have any opinion on this proposition.” Such responses can not be mapped to any survey option or Likert scale. We classify these responses as unclassifiable and label them as “00”.
To validate this process, the first two authors manually reviewed a random sample of $100$ forced open-ended and $100$ fully unconstrained responses, extracting stances and comparing them to the model-generated stances for the corresponding questions. This cross-matching achieved an overall accuracy of $0.957$.
@@ 3.3 Evaluation Metrics
Here, $N$ is the total number of questions, $y_{i}$ represents the ground truth for the $i$-th question, and $y_{i}^{\prime}$ is the model’s response. The indicator function ${1}{\{y_{i}=y_{i}^{\prime}\}}$ equals $1$ if $y_{i}=y_{i}^{\prime}$ and 00 otherwise.
@@
$\epsilon_{i}=\begin{cases}\frac{|y_{i}^{\prime}-y_{i}|}{q_{i}-1}&\text{if the question is ordinal},\\ 1-\text{reward}&\text{if the question is categorical}.\end{cases}$
Here, $q_{i}$ denotes the number of options for the $i$-th question (for ordinal questions). For categorical questions, the reward is proportional to the number of matching elements between the ground truth and the model response. This approach ensures a nuanced evaluation of model performance across different question types.
@@ 4.1 World Values Survey (WVS)
Traditional closed-style classification often fails to capture the full extent of cultural alignment in LLMs under less constrained conditions. As illustrated in Table 1, the standard closed-style MCQ-based approach achieves the highest alignment in only approximately $20\%$ of cases in hard alignment metric, and an even lower $13.3\%$ in soft alignment metric when prompting in native language. This performance improves slightly to $33.3\%$ for hard alignment when the prompts are provided in English. On the other hand, the two less constrained settings dominate in the majority of cases. With the exception of a few outliers, specifically for Bangladesh, all models demonstrate their strongest alignment performance in these less constrained settings, while some also perform notably well in the reverse order scenario.
@@
Table 2: Unclassifiable Outputs for Different Probing Methods across the Countries Studied
@@
(a) GPT-4o
(b) GPT-4
(c) Llama 3.3
(d) Mistral Large 2
(e) DeepSeek R1
Figure 4: Comparison of cross-cultural correlation per value among the four probing methods for all models. The plot uses Paul Tol’s Colorblind Palette to ensure vision accessibility. Numeric results are presented in Table 11.
@@
The unclassifiable outputs in less constrained settings provide a wealth of information about socio-cultural characteristics. As illustrated in Table 2, the models frequently refused to answer certain questions in the two less constrained probing methods (Forced Open-Ended and Fully Unconstrained) for Germany and the USA. To better understand this phenomenon, we conduct a case study to investigate which questions (full question list in Table 5) are predominantly avoided by the models. Our findings indicate that questions related to the importance of God and the acceptability of abortion in society are most often met with non-specific responses from the models. For instance, when asked about the importance of God, the models respond with statements such as ’One’s relationship with God or a higher power is a private matter, and not something that should be imposed on others’. Similarly, when questioned on the acceptability of abortion, responses include statements like ’This is a decision that should be left up to the individual woman, in consultation with her doctor and loved ones’. We hypothesize that these responses reflect an accurate representation of the societal values of Germany and the USA. As discussed in Section 2.4, both countries are positioned within the secular-rational and self-expression dimension of the Inglehart-Welzel Cultural Map, which is characterized by societies that prioritize subjective well-being, civic activism, and self-expression. These traits are particularly prominent in postindustrial societies with high levels of existential security and individual autonomy (Inglehart, 2005). The nuanced responses observed in these open-ended settings align closely with the defining characteristics of such societies. Notably, these subtleties are entirely unattainable in the closed-style survey settings. The contrast is illustrated in Figure 3 for Germany using two language models: Llama 3.3 and GPT-4. This difference becomes even more apparent through the thematic analysis of WVS responses detailed in Appendix D.3.
@@
Responses exhibit significant sensitivity to the order of options in both MCQ and Likert scale-based questions from the WVS, with notable improvements (ranging from $3\%-15\%$) in soft metric in majority of the cases when the order of options is reversed. This sensitivity highlights the positional bias inherent in LLMs (Pezeshkpour and Hruschka, 2023), where the placement of options—such as in ascending or descending order—can significantly influence their preferences. We argue that in a nuanced and critical domain like cultural alignment, this positional fragility represents a fundamental flaw in the evaluation framework, as it undermines the reliability of the results when LLMs are highly susceptible to such superficial variations.
@@ 4.2 Hofstede Cultural Dimension
Similar to the WVS results presented in Section 4.1, less constrained settings achieve a higher level of positive correlation with statistical significance in the Hofstede cultural survey. As illustrated in Table 3, $66.67\%$ of the cases of cross-value correlation per country are dominated by the two less constrained probing methods: forced open-ended and fully unconstrained. Notably, all statistically significant correlations ($p<=0.05$) are exclusively achieved by these two methods. We further evaluate cross-cultural correlations per value across the five models. The findings, as depicted in Figure 4 and Table 11, reveal that while the two closed-style probing methods achieve statistically significant positive correlations in $33.34\%$ of cases across all six dimensions, the two unconstrained settings demonstrate statistically significant results in $58.34\%$ of cases across all dimensions for all models.
One notable observation from Table 3 is that, when prompted in Bengali, the two unconstrained generation settings exhibited suboptimal performance, frequently yielding negative correlations-a trend markedly distinct from the results observed for the other two cultures examined in this study, Germany and the USA. Upon closer examination of the responses generated by these unconstrained methods, it becomes evident that the models often fail to provide detailed descriptions or substantive reasoning to support their answers in Bengali. We hypothesize that this underperformance is due to Bengali being a resource-scarce language in the field of natural language processing (NLP), especially when compared to languages like English and German. This disparity arises from limited availability of datasets, underdeveloped NLP tools, and a lack of research focus on Bengali (Joshi et al., 2020). Consequently, the unconstrained generation methods struggle to perform effectively for the Bangladeshi cultural context.
@@ 5 Reconsidering Cultural Alignment
Survey-based closed-style questionnaires are insufficient for accurately evaluating cultural alignment. Anthropological and social science perspectives emphasize the intricate nature of culture, which extends far beyond observable behaviors to encompass the lived experiences of individuals within a society (Geertz, 2017), which cannot be adequately captured through a closed-style multiple-choice question (MCQ) format. Such approach restricts language models by preventing them from expressing uncertainty or refusing to answer—behaviors that can offer valuable insights into societal structures (Urman and Makhortykh, 2023), as discussed in Section 4.1 of our study. Instead, it risks creating an artificial sense of alignment through multiple-choice (MCQ) responses, a phenomenon well-documented in prior literature. A more comprehensive conceptualization of culture views it as an amalgamation of demographic and semantic proxies, each contributing to the complex tapestry of societal norms, values, and behaviors (Adilazuarda et al., 2024), with different evaluation frameworks tailored to assess specific proxies. In this regard, the studies by Naous et al. (2023); Rao et al. (2024) provide valuable examples by developing specialized frameworks to evaluate the adaptability of large language models (LLMs) across various cultural proxies, such as food, regional distinctions, and social values. Based on these insights, we recommend that future research should move beyond the exclusive reliance on closed-style evaluations and adopt more flexible, framework-based methodologies. This approach would allow for a more holistic and accurate assessment of language models’ ability to understand and navigate the complexities of diverse cultural landscapes.
@@
The assertion that language can serve as a reliable proxy for culture is a contentious and potentially flawed approach in academic research. While previous studies have used language as a representative indicator for cultural alignment and understanding (AlKhamissi et al., 2024; Masoud et al., 2023), this methodology oversimplifies the complex relationship between language and culture. The acquisition of a language associated with a particular culture is often a relatively straightforward process that can be accomplished within a relatively short timeframe. In contrast, the internalization and adoption of cultural values is a far more intricate and time-consuming process, often spanning generations or even centuries to become deeply ingrained within a society (Smolicz, 1981). In their study, Havaldar et al. (2023) demonstrate that multilingual models fail to fully capture cultural variations tied to emotions and predominantly reflect the cultural values of Western societies. The ability of a language model to process or generate content in a specific language should not be equated with a comprehensive grasp of the associated cultural values and emotional expressions. This critique is further exemplified by the findings discussed in section 4.2, where the poor performance of unconstrained generation methods in Bengali highlights the practical limitations of NLP models in resource-scarce languages. Together, these observations reinforce the need for both technical advancements in low-resource languages and a conceptual shift beyond relying on language as a sole proxy for culture.
@@
Surveys like WVS and Hofstede can serve as valuable resources of knowledge or fine-tuning data for integrating cultural differences in LLMs , rather than being solely used for evaluating cultural alignment. A notable example of this approach is CultureLLM (Li et al., 2024a), where the authors utilize augmented data derived from WVS survey responses to fine-tune LLMs. After fine-tuning, they evaluate the models’ cultural adaptability by testing them on culture-related public datasets in specific languages, observing measurable improvements. This demonstrates the potential of integrating survey data not only for testing, but also for improving the cultural sensitivity and adaptability of LLMs.

### v3

@@ 1 Introduction
To fully unlock the potential of artificial intelligence and achieve the vision of “AI for All”, it is essential to design and develop the Large Language Models (LLMs) with a focus on inclusivity and relevance across diverse societies and cultures (Adilazuarda et al., 2024). However, the recent advancements in language models have faced significant criticism for predominantly reflecting Western perspectives (Durmus et al., 2023) and displaying pronounced biases toward Anglocentric or American cultural norms (Johnson et al., 2022; Dwivedi et al., 2023). Such biases pose serious risks, including the stereotyping and alienation of users from underrepresented communities due to the lack of cultural sensitivity and understanding (Cao et al., 2022). Addressing this issue begins with identifying and understanding the problem. This raises a critical question: to what extent do current LLMs align with the diverse cultural contexts worldwide, and how can this alignment be systematically evaluated?
@@
  • We demonstrate that closed-style questions alone are insufficient for a comprehensive evaluation of LLMs for cultural alignment, as unconstrained prompts often yield richer and more insightful responses and exhibit stronger alignment across cultures.
@@ 2.1 World Values Survey
The World Values Survey (WVS) (Haerpfer et al., 2024) is a global research initiative consisting of approximately 250 questions, organized into 14 thematic sections covering topics such as social values, stereotypes, societal well-being, and economic values. Political scientists Ronald Inglehart and Christian Welzel analyzed the WVS data and identified two key dimensions of cross-cultural variation: Traditional versus Secular-rational values and Survival versus Self-expression values. These dimensions are derived through a factor analysis of $10$ indicators presented in Table 6. Although these indicators represent only a subset of beliefs and values, they effectively capture critical aspects of cultural variation. We adopt these $10$ closed-style indicators/questions as our first probing corpus.
@@ 2.3 Probing Methods
We concatenate Anthropological Prompting (AlKhamissi et al., 2024) before each probing, which grounds questions in anthropological contexts by guiding the model to think as if actively participating in the method.The authors identify six demographic dimensions and empirically find that optimal cultural alignment occurs with the following specific attributes: Region: Country-Specific, Sex: Male, Age: <50<50, Social Class: Upper/Lower Middle Class, Education Level: Higher, and Marital Status: Married.We adopt this configuration to allow our models the best opportunity to align with the cultural contexts under study. Our complete probing methodology is illustrated in Figure 2, using one example question from both the WVS and Hofstede surveys.
@@
2.4 Country & Language Selection
To ensure a robust analysis of cross-cultural differences, we select countries based on their positions on the Inglehart-Welzel World Cultural Map 222www.worldvaluessurvey.org/WVSNewsShow.jsp?ID=467, which is structured around two key dimensions: Traditional versus Secular-rational values and Survival versus Self-expression values. This method allows us to capture contrasting cultural orientations while ensuring the availability of ground truth data for validation. Additionally, we prompt language models in both English and the native languages of the selected countries to enhance cultural alignment and correlate results with survey findings (Cao et al., 2023a). Based on these criteria, we select Bangladesh, Germany, and the USA. Bangladesh represents a society deeply rooted in Traditional and Survival values, while Germany aligns strongly with Secular-rational and Self-expression values. To provide a reference point, we include the USA, as it is frequently used in cross-cultural studies and because large language models (LLMs) often exhibit bias toward American culture (Johnson et al., 2022). The USA also occupies the Secular-rational and Self-expression dimension, similar to Germany on the Inglehart-Welzel map. For Bangladesh and Germany, we source survey questions in Bengali and German from the official survey websites, which provide translations in numerous languages. We assume English to be the native language of the USA in our study. Finally, we extend our experiments to the Philippines which is culturally proximate to Bangladesh on the Inglehart–Welzel map, to test additional hypotheses later in the paper.
@@ 3.1 Models
Table 1: Cultural alignment comparison using World Values Survey (WVS) when prompted in English and Native languages for Germany, Bangladesh, and the USA. Scores(%) are in Hard / Soft alignment metrics for each model using four probing method: FC (Forced Closed-Style), FR (Forced Reverse Order), FO (Forced Open-ended), FU (Fully Unconstrained). Higher percentages indicate stronger cultural alignment, with the boldfaced values representing the highest alignment scores per country across the four probing methods for each model. The columns for the USA remain the same across languages assuming English to be the native language.
@@ 3.2 Evaluation
For the first two prompting methods, forced closed-style and forced reverse order, model responses are obtained in numerical form, consistent with traditional survey options. We use a straightforward Python script to map the responses from the reverse order questions back to the original survey option for evaluation. For the unconstrained settings (forced open-ended and fully unconstrained), we follow the approach of Bravansky et al. (2025) to map free-form responses to the most appropriate survey options. Specifically, we use GPT-4o to analyze these responses and determine the model’s stance using the prompt shown in Figure 6. In many cases, LLMs respond to certain questions with statements like, “As an AI, I do not have any opinion on this proposition.” Such responses can not be mapped to any survey option or Likert scale. We classify these responses as unclassifiable and label them as “0”.
To validate this mapping procedure, we adopt two cautionary steps:
  1. 1.
    The first two authors manually reviewed a random sample of $100$ forced open-ended and $100$ fully unconstrained responses, comparing extracted stances with the model-generated ones. This cross-check yielded an overall accuracy of $0.957$.
  2. 2.
    To mitigate single-LLM bias, we additionally annotate $100$ unconstrained responses using Claude Sonnet 3.7, an independent model not used elsewhere in this study. The comparison between GPT-4o and Claude Sonnet 3.7 achieved a Cohen’s kappa of $0.81$, indicating almost perfect agreement.
@@ 3.3 Evaluation Metrics
Here, $N$ is the total number of questions, $y_{i}$ represents the ground truth for the $i$-th question, and $y_{i}^{\prime}$ is the model’s response. The indicator function ${1}{\{y_{i}=y_{i}^{\prime}\}}$ equals $1$ if $y_{i}=y_{i}^{\prime}$ and 0 otherwise.
@@
[…]
Here, $q_{i}$ denotes the number of options for the $i$-th question (for ordinal questions). For categorical questions, the reward is proportional to the number of matching elements between the ground truth and the model response. This approach ensures a nuanced evaluation of model performance across different question types. Both soft and hard metric scores are reported as percentages, where higher values indicate stronger alignment with the respective culture.
@@ 4.1 World Values Survey (WVS)
Traditional closed-style classification often fails to capture the full extent of cultural alignment in LLMs under less constrained conditions. As illustrated in Table 1, the standard closed-style MCQ-based approach achieves the highest alignment in only approximately $20\%$ of cases in the hard alignment metric and an even lower $13.3\%$ in the soft alignment metric when prompting in the native language. This performance improves slightly to $33.3\%$ for hard alignment when the prompts are provided in English. On the other hand, the two less constrained settings dominate in the majority cases. With the exception of a few outliers, specifically for Bangladesh, all models demonstrate their strongest alignment performance in these less constrained settings, while some also perform notably well in the reverse order scenario.
@@
Table 2: Unclassifiable outputs across different probing methods. All cases arise from unconstrained settings.
@@
Figure 4: Comparison of cross-cultural correlation per value across the four probing methods for all models. Color intensity encodes the correlation strength (deep blue = strong positive, deep red= strong negative, white = near-zero). Values closer to 1 indicate stronger alignment with expected cultural dimensions, and starred values (striped cells) highlight the values with statistical significance. Notably, the unconstrained settings achieve the highest proportion of statistically significant positive correlations, highlighting its stronger alignment performance across models and dimensions.
@@
The unclassifiable outputs in less constrained settings provide a wealth of information on socio-cultural characteristics. As illustrated in Table 2, the models frequently refused to answer certain questions in the two less constrained probing methods (Forced Open-Ended and Fully Unconstrained) for Germany and the USA. To better understand this phenomenon, we conduct a case study to investigate which questions (full question list in Table 6) are predominantly avoided by the models. Our findings indicate that questions related to the importance of God and the acceptability of abortion in society are most often met with non-specific responses from the models. For instance, when asked about the importance of God, the models respond with statements such as ‘One’s relationship with God or a higher power is a private matter, and not something that should be imposed on others’. Similarly, when questioned on the acceptability of abortion, responses include statements like ‘This is a decision that should be left up to the individual woman, in consultation with her doctor and loved ones’. We hypothesize that these responses reflect an accurate representation of the societal values of Germany and the USA. As discussed in Section 2.4, both countries are positioned within the secular-rational and self-expression dimension of the Inglehart-Welzel Cultural Map, which is characterized by societies that prioritize subjective well-being, civic activism, and self-expression. These traits are particularly prominent in postindustrial societies with high levels of existential security and individual autonomy (Inglehart, 2005). The nuanced responses observed in these open-ended settings align closely with the defining characteristics of such societies. To validate this hypothesis, we extend our investigation to the Philippines, a nation positioned in close cultural proximity to Bangladesh on the Inglehart-Welzel map to determine whether unclassifiable outputs represent a generalizable phenomenon or are specific to certain cultural contexts like Bangladesh. We can confirm that neither the Philippines nor Bangladesh yield any unclassifiable output, strengthening our hypothesis that such responses are distinctive markers of secular-rational, self-expression-oriented societies rather than universal model behaviors. Notably, these subtleties are entirely unattainable in the closed-style settings. The contrast is illustrated in Figure 3 for Germany and the USA, where Llama 3.3 achieves closer alignment under unconstrained probing compared to closed-style settings. In contrast, for Bangladesh, constrained methods perform on par with unconstrained ones. The WVS results in Figure 9 likewise show that constrained and unconstrained settings achieve comparable cultural alignment for both Bangladesh and the Philippines.
@@
Responses exhibit significant sensitivity to the order of options in both MCQ and Likert scale-based questions from the WVS, with notable improvements (ranging from $3\%-15\%$) in soft metric in the majority of the cases when the order of options is reversed. This sensitivity highlights the positional bias inherent in LLMs (Pezeshkpour and Hruschka, 2023), where the placement of options, such as in ascending or descending order, can significantly influence their preferences. We argue that in a nuanced and critical domain like cultural alignment, this positional fragility represents a fundamental flaw in the evaluation framework, as it undermines the reliability of the results when LLMs are highly susceptible to such superficial variations.
@@ 4.2 Hofstede Cultural Dimension
Similar to the WVS results presented in Section 4.1, less constrained settings achieve a higher level of positive correlation with statistical significance in the Hofstede cultural survey. As illustrated in Table 3, $66.67\%$ of the cases of cross-value correlation per country are dominated by the two less constrained probing methods: forced open-ended and fully unconstrained. Notably, all statistically significant correlations ($p<=0.05$) are exclusively achieved by these two methods. We further evaluate cross-cultural correlations per value across the five models. The findings, as depicted in Figure 4 and Table 12, reveal that while the two closed-style probing methods achieve statistically significant positive correlations in $33.34\%$ of cases across all six dimensions, the two unconstrained settings demonstrate statistically significant results in $58.34\%$ of the cases across all dimensions for all models.
Model Philippines
FC FR FO FU
GPT-4o 0.17 / 0.17 0.12 / 0.17 0.52 / 0.06 0.37 / 0.31
Llama 3.3 0.12 / 0.41 0.37 / 0.43 0.29 / 0.20 0.20 / -0.03
Mistral L2 0.26 / 0.06 -0.06 / 0.39 0.46 / 0.15 0.37 / -0.20
DeepSeek-R1 0.51 / 0.06 0.23 / 0.17 0.58 / 0.64 0.83* / 0.83*
Table 4: Cross-value correlation scores for the Philippines on the Hofstede Cultural Dimension. Scores for each probing method are presented as Prompted in English / Prompted in Filipino. Statistically significant ($p<=0.05$) scores are marked with *.
However, correlations exhibit limited statistical significance in the unconstrained setting in general, and even weaker significance in closed-style settings, a phenomenon also observed in previous work (Arora et al., 2023), which reports similarly low significance in Hofstede dimensions. One notable observation from Table 3 is that, when prompted in Bengali, the two unconstrained generation settings exhibit suboptimal performance, frequently yielding negative correlations, which is markedly distinct from the results observed for the other two cultures examined in this study, Germany and the USA. Upon closer examination of the responses generated by these unconstrained methods, it becomes evident that the models often fail to provide detailed descriptions or substantive reasoning to support their answers in Bengali. We initially hypothesize that this performance degradation stems from Bengali’s status as a low-resource language within the NLP ecosystem, particularly when contrasted with well-resourced languages like English. To rigorously test this hypothesis, we expand our investigation to include the Philippines as mentioned in Section 4.1, another nation whose primary language (Filipino) occupies a similar low-resource position in NLP. As Table 4 demonstrates, Filipino prompts similarly yield negative correlations, predominantly in unconstrained settings which corroborates our initial observations. Unexpectedly, this pattern is also manifested in German unconstrained settings, despite German’s classification as a relatively well-resourced language in NLP.
This unexpected finding points to a more fundamental limitation than simple resource availability. The observed performance disparities likely reflect a complex interplay of factors: limited training data diversity, insufficient cross-linguistic representation in model architectures, and the overwhelming research emphasis on English-language benchmarks (Joshi et al., 2020). These systemic biases create cascading effects wherein models default to English-centric reasoning patterns, resulting in culturally incongruent or semantically impoverished responses during text generation in non-English linguistic contexts, even for languages with moderate resource availability.
5 Annotation Divergence: Human vs. LLM
As described in Section 3.2, we validate the annotation of open-ended responses by GPT-4o through a manual review conducted independently by the first two authors. To ensure neutrality, we adopt a blinding protocol in which reviewers are unaware of:
  • the source of the response (originating LLM),
  • the adjudication of annotators’ ratings prior to the final reconciliation.
Reviewers were instructed to minimize subjective influence and maintain impartiality throughout the process. Any discrepancies that emerged after the initial annotation are resolved through consensus.
We further examine cases where manual and LLM annotations diverge for open-ended responses. For WVS questions, disagreements primarily occur on questions concerning the importance of God or the justifiability of homosexuality or abortion, which also yielded many unclassifiable outputs (see Section 4.1). We attribute this to the inherently subjective nature of these questions and the challenge of mapping nuanced responses to a continuous 111010 scale, where $1$ indicates no agreement and $10$ indicates full agreement. While LLMs readily assign a score within this range, human annotators often find it difficult to map ambiguous or hedged responses to a deterministic value.
For Hofstede Cultural Dimension, although many questions are also subjective (e.g., “How proud are you to be a citizen of your country? (1) very proud (2) fairly proud (3) somewhat proud (4) not very proud (5) not proud at all.”), responses are mapped to a 1155 scale. This reduces the granularity of possible values, and thereby limits the scope for divergence between manual and LLM annotations.
@@ 6 Reconsidering Cultural Alignment
Survey-based closed-style questionnaires are inadequate to accurately evaluate cultural alignment in LLMs. Anthropological and social science perspectives emphasize the intricate nature of culture, which extends far beyond observable behaviors to encompass lived experiences of individuals within a society (Geertz, 2017), which cannot be adequately captured through a closed-style multiple-choice question (MCQ) format. Such approach restricts language models by preventing them from expressing uncertainty or refusing to answer, behaviors that can offer valuable insights into societal structures (Urman and Makhortykh, 2023), as discussed in Section 4.1 of our study. Instead, it risks creating an artificial sense of alignment through multiple-choice (MCQ) responses, a phenomenon well-documented in prior literature. A more comprehensive conceptualization of culture views it as an amalgamation of demographic and semantic proxies, each contributing to the complex tapestry of societal norms, values, and behaviors (Adilazuarda et al., 2024), with different evaluation frameworks tailored to assess specific proxies. In this regard, the studies by Naous et al. (2023); Rao et al. (2024) provide valuable examples by developing specialized frameworks to evaluate the adaptability of large language models (LLMs) across various cultural proxies, such as food, regional distinctions, and social values. Thus, a practical approach for benchmarking the cultural alignment of LLMs is to begin with demographic and cultural proxies (Adilazuarda et al., 2024) and evaluate LLMs’ understanding around these proxies in order to have culturally aware benchmarks. This approach would allow for a more holistic and accurate assessment of language models’ ability to understand the complexities of diverse cultural landscapes.
@@
The assertion that language can serve as a reliable proxy for culture is a contentious and potentially flawed approach in academic research. While previous studies have used language as a representative indicator for cultural alignment and understanding (AlKhamissi et al., 2024; Masoud et al., 2023), this methodology oversimplifies the complex relationship between language and culture. The acquisition of a language associated with a particular culture is often a relatively straightforward process that can be accomplished within a relatively short time-frame. In contrast, the internalization and adoption of cultural values is a far more intricate and time-consuming process, often spanning generations or even centuries to become deeply ingrained within a society (Smolicz, 1981). Havaldar et al. (2023) demonstrate that multilingual models fail to fully capture the cultural variations tied to emotions and predominantly reflect the cultural values of Western societies in their study. The ability of a language model to process or generate content in a specific language should not be equated with a comprehensive grasp of the associated cultural values and emotional expressions. This critique is evident in the findings discussed in Section 4.2, where the poor performance of unconstrained generation methods in Bengali and Filipino highlights the practical limitations of NLP models in resource-scarce languages. The challenge extends beyond resource scarcity; even well-resourced languages show cultural misalignment, compounded by the multifaceted nature of cultural identity, spanning religion, nationality, ethnicity, and regional variation, which remains severely underrepresented in NLP benchmarks (Das et al., 2023). Together, these observations reinforce the need for both technical advancements in low-resource languages and a conceptual shift beyond relying on language as the sole proxy for culture.
@@
Surveys like WVS and Hofstede can serve as valuable resources of knowledge or fine-tuning data for integrating cultural differences in LLMs , rather than being solely used for evaluating cultural alignment. A notable example of this approach is CultureLLM (Li et al., 2024a), where the authors utilize augmented data derived from WVS survey responses to fine-tune LLMs. After fine-tuning, they evaluate the models’ cultural adaptability by testing them on culture-related downstream tasks (e.g. stance detection, offensive language detection, etc) in specific languages and observe measurable improvements. This demonstrates the potential of integrating survey data not only for testing, but also for improving the cultural sensitivity and adaptability of LLMs.